# Compressed Feature Quality Assessment: Dataset and Baselines

### v1

@@ Abstract.
The widespread deployment of large models in resource-constrained environments has underscored the need for efficient transmission of intermediate feature representations. In this context, feature coding, which compresses features into compact bitstreams, becomes a critical component for scenarios involving feature transmission, storage, and reuse. However, this compression process introduces inherent semantic degradation that is notoriously difficult to quantify with traditional metrics. To address this, this paper introduces the research problem of Compressed Feature Quality Assessment (CFQA), which seeks to evaluate the semantic fidelity of compressed features. To advance CFQA research, we propose the first benchmark dataset, comprising 300 original features and 12000 compressed features derived from three vision tasks and four feature codecs. Task-specific performance drops are provided as true semantic distortion for the evaluation of CFQA metrics. We assess the performance of three widely used metrics (MSE, cosine similarity, and Centered Kernel Alignment) in capturing semantic degradation. The results underscore the representativeness of the dataset and highlight the need for more refined metrics capable of addressing the nuances of semantic distortion in compressed features. To facilitate the ongoing development of CFQA research, we release the dataset and all accompanying source code at https://github.com/chansongoal/Compressed-Feature-Quality-Assessment. This contribution aims to advance the field and provide a foundational resource for the community to explore CFQA.
Compressed Feature Quality Assessment, Coding for Machines, Feature Coding
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Multimedia databasesccs: Computing methodologies Image compressionccs: Human-centered computing Ubiquitous and mobile computing design and evaluation methods
@@ 1. Introduction
The rapid deployment of large foundation models (e.g., DINOv2, LLaMA3) across distributed and resource-constrained environments has created a growing need to transmit intermediate features rather than raw signals. Feature coding, which compresses these intermediate representations, plays a vital role in enabling scalable, privacy-preserving, and efficient systems. Unlike traditional image or video coding that prioritizes perceptual quality, feature coding targets the preservation of task-relevant semantics embedded in deep representations.
@@
However, feature coding inevitably introduces semantic degradation: a loss of discriminative information that can compromise downstream performance. This degradation is fundamentally different from pixel-level distortions and often cannot be captured by conventional distortion metrics such as MSE or PSNR. While task accuracy is a more reliable indicator of semantic quality, it is impractical for deployment: downstream tasks may be inaccessible, costly to run, or unavailable at compression time. These limitations highlight an urgent and underexplored problem: Compressed Feature Quality Assessment (CFQA) — how can we estimate the semantic quality of compressed features without relying on downstream inference?
@@
Solving CFQA presents several challenges. First, there is no standard dataset that provides compressed features with corresponding task performance across multiple tasks and codecs. Second, existing similarity metrics lack validation for high-dimensional features and often fail to generalize across tasks. Third, there is no unified protocol for evaluating whether a given metric faithfully reflects semantic degradation.
@@
To bridge this gap, we take the first step toward systematic study of CFQA. In contrast to prior works that rely on task supervision or task-specific codecs, we treat CFQA as a standalone problem and aim to benchmark its core components: datasets, metrics, and evaluation protocols.
We make the following contributions.
@@
  • Multi-Task Benchmark Dataset: We construct the first CFQA dataset, consisting of 300 original features and 12,000 compressed features from three vision tasks (classification, segmentation, depth estimation) and four representative codecs (including handcrafted and learned models). The dataset enables quantitative analysis of semantic degradation across tasks, bitrates, and codecs.
@@
  • Truth Semantic Distortion Derivation: For each compressed feature, we provide task-specific semantic distortion labels (rank change, mIoU drop, RMSE difference) computed by comparing task head outputs using original and compressed features. These labels serve as the ground truth for training and evaluating quality metrics.
@@
  • Existing Metric Evaluation: We evaluate three representative signal-based metrics(MSE, cosine similarity, and Centered Kernel Alignment (CKA)) and analyze their sensitivity to compression and their correlation with task-specific degradation. Our analysis reveals their varying effectiveness, stability, and failure cases, providing critical insights into their applicability for semantic quality prediction.
@@
We hope this work lays the foundation for future CFQA research and bridges the gap between low-level compression and high-level semantic utility. By providing a standardized benchmark and analysis pipeline, our work encourages the development of lightweight, generalizable, and task-agnostic quality metrics that support real-world deployment of compressed features.
@@ 2.1. Feature Coding
Feature coding has received growing attention (Zhang et al., 2021; Kim et al., 2023; Misra et al., 2022; Alvar and Bajić, 2019, 2020; Guo et al., 2023; WG2, 2023; Yan et al., 2021; Choi and Bajić, 2018; Chen et al., 2020b, a; Gao et al., 2024a; Zhu et al., 2024; Gao et al., 2025; Ma et al., 2024) in the cloud-edge collaborative intelligence scenarios at early stage. Recently, large model feature coding has attracted increasing interest (Oquab et al., 2023; Guo et al., 2025). These work extend feature coding to the scenarios where features are transmitted, stored, and reused.
@@ 2.2. Semantic Distortion Measurement
Task Source Num. of Org. Feat. Propocessing Feature Codecs Num of Comp. Feat. Feature Shape GT Distortion
Cls ImageNet 100 Resize HM, VTM Multi-task Hyperprior Task-specific Hyperprior 4000 257$\times$1536 Rank
Seg VOC 2012 100 Flip and Crop 4000 2$\times$1370$\times$1536 mIoU difference
Dpt NYUv2 100 Flip 4000 2$\times$4$\times$161$\times$1536 RMSE difference
@@ 3.1. Problem Formulation
Given a pretrained model $\mathcal{M}$, we denote its extracted intermediate feature from an input $x\in\mathcal{X}$ as $\mathbf{f}=\mathcal{M}(x)\in\mathbb{R}^{d}$, where $d$ is the feature dimension. A feature codec $\mathcal{C}$ encodes $\mathbf{f}$ into a compact bitstream and decodes a compressed representation $\hat{\mathbf{f}}=\mathcal{C}^{-1}(\mathcal{C}(\mathbf{f}))$. Our objective is to evaluate how much semantic information is preserved in $\hat{\mathbf{f}}$ with respect to $\mathbf{f}$, without accessing downstream task ground-truth or executing inference.
We define this evaluation task as Feature Quality Assessment (FQA): Given a pair of original and compressed features $(\mathbf{f},\hat{\mathbf{f}})$, estimate the semantic quality $q\in\mathbb{R}$ of the compressed feature, such that $q$ strongly correlates with its performance on downstream tasks.
Assume a downstream task $\mathcal{T}$ with a head network $h_{\mathcal{T}}$, producing an output $y=h_{\mathcal{T}}(\cdot)$. Let the task-specific performance metric be denoted as $\mathcal{A}_{\mathcal{T}}(\cdot)$. When using compressed features, the task performance becomes: $s=\mathcal{A}_{\mathcal{T}}(h_{\mathcal{T}}(\hat{\mathbf{f}}))$. This value $s\in\mathbb{R}$ reflects the ground-truth semantic utility of the compressed feature. However, computing $s$ requires task execution and labels, which may be unavailable or expensive. The goal of FQA is to estimate a score $q=Q(\mathbf{f},\hat{\mathbf{f}})$, where $Q(\cdot)$ is a task-agnostic quality metric, such that $q\approx s$ in terms of correlation across samples.
@@ 3.2. Application Scenarios
We illustrate two typical application scenarios of CFQA in Fig. 1.
First, CFQA can be directly integrated into the codec training process as a supervisory signal to align compression objectives with downstream task performance. Recent studies (Singh et al., 2020; Gao et al., 2024a) show that task-aware feature coding benefits significantly from semantic-aware supervision. CFQA offers such guidance without requiring end-to-end task labels. As shown on the left of Fig. 1, both the original and reconstructed features are input to the CFQA module, which estimates the semantic distortion. The distortion score is then used to generate gradients for optimizing the encoder and decoder. With a well-designed CFQA metric, the resulting codec learns to preserve semantic content even under low-bitrate constraints.
Second, CFQA is essential in edge-cloud collaborative systems, where features are extracted at the edge and transmitted to the cloud for inference. Since the downstream model resides on the cloud side, semantic distortion cannot be directly measured at the edge. Here, CFQA is used as a proxy to estimate task-relevant degradation. As illustrated on the right of Fig. 1, the compressed feature is first evaluated by the CFQA module. If it is judged to be of high quality, the bitstream is transmitted to the cloud. Otherwise, the edge device re-encodes the feature at a higher bitrate to better preserve semantic information. This strategy ensures the reliability of transmitted features while reducing unnecessary bandwidth consumption.
Although our work focuses on semantic degradation due to compression, we emphasize that the value of CFQA extends beyond codec benchmarking: it serves as a crucial component in a variety of systems where features are extracted, compressed, transmitted, cached, or reused.
@@ 4.1. Overview
Our proposed dataset is designed to support the analysis of semantic distortion introduced by lossy compression. It includes three tasks, 300 original features, and 12000 compressed features from 4 kinds of feature codecs. The feature cover diverse image processing methods and multiple feature extraction strategies. The abstract information of the dataset is presented in Table 1.
@@ 4.2. Model and Task Selection
Since feature coding still focus on visual signals, we start CFQA by visual features. We choose DINOv2 (Oquab et al., 2023) as the backbone model for feature extraction due to its strong generalization ability and wide adoption in general-purpose vision tasks. To cover a broad range of semantic complexity, we select three widely studied vision tasks: image classification (Cls), semantic segmentation (Seg), and depth estimation (Dpt). These tasks span coarse-to-fine semantic understanding: Cls focuses on image-level category prediction, Seg introduces spatial semantics with class-wise alignment at the pixel level, and Dpt requires detailed geometry prediction.
By including these three tasks, we aim to ensure that semantic distortion is evaluated across multiple levels of abstraction, which is essential for assessing the general applicability and sensitivity of CFQA metrics.
@@ 4.3. Source Data Collection
To ensure semantic diversity, we select 100 representative samples for each task. For Cls, we sample 100 ImageNet (Deng et al., 2009) images from 100 distinct categories, each correctly predicted by the DINOv2 classifier. For Seg, we sample 100 images from the Pascal VOC 2012 (Everingham et al., 2010) validation set, covering all 20 semantic categories. For Dpt, we sample 100 images from the NYUv2 (Nathan Silberman and Fergus, 2012) dataset, spanning all 16 scenes.
This sampling ensures feature diversity while maintaining dataset size suitable for controlled evaluation.
@@ 4.4. Original Feature Collection
The various image preprocessing and split points simulate real-world input variation and improve generalizability.
@@ 4.5. Compressed Feature Collection
To simulate different types and strengths of semantic degradation, we compress original features through four codecs. All original features are flattened into 2D arrays before encoding, and compressed features are reshaped to original formats after decoding.
@@
We uniformly quantize original features to [0, 1023], followed by two handcrafted codecs: HM Intra coding (configured with encoder_intra_main_rext.cfg) and VTM Intra coding: (configured with encoder_intra_vtm.cfg). For both codec, we use the YUV-400 format for feature coding and the quantization parameters set to $\in{2,4,6,\ldots,20}$ to simulate different bitrate levels and distortion strengths.
@@
We adopt the Hyperprior model (Ballé et al., 2018), which serves as the milestone for learned feature coding. Hyperprior models offer a good balance between modeling capacity and computational efficiency. Most existing compression methods follow the Hyperprior architecture. Therefore, it can simulate the semantic distortion introduced by learning-based methods.
To analyze optimization effects, we train two variants: Task-Specific: The codec is trained independently on features extracted from a specific task. Multi-Task: The codec is trained on features extracted from the three tasks. All codecs are trained using the training strategies proposed in (Gao et al., 2024c), and the lambda information can be found in our released models.
4.6. True Semantic Distortion Collection
To evaluate the effectiveness of CFQA metrics, we need to compare them with true semantic distortion. We define the true semantic distortion distortion as the task performance difference between the compressed featrure $\hat{\mathbf{f}}$ and the original feature $\mathbf{f}$.
For Cls, we compute the rank of the ground-truth label in the prediction logits derived from $\hat{\mathbf{f}}$. Higher rank indicates greater semantic degradation. The rank value of the original feature is 1. For Seg, we compute the mIoU difference between the segmentation masks predicted from $\hat{\mathbf{f}}$ and $\mathbf{f}$. For Dpt, we compute the RMSE difference between the depth maps predicted from $\hat{\mathbf{f}}$ and $\mathbf{f}$.
These task-specific scores serve as ground-truth semantic distortion for evaluating the effectiveness of feature quality metrics.
@@ 4.6. True Semantic Distortion Collection
Multi-Task-Trained Hyperprior Task-Specific-Trained Hyperprior
@@
Hyperprior (Multi-Task) Cls -0.0281 0.4707 -0.8907 -0.7165 -0.3277 -0.6593
@@
Hyperprior (Task-Specific) Cls -0.0220 0.1303 -0.5486 -0.6084 -0.2622 -0.3365

### v2

@@ Abstract.
The widespread deployment of large models in resource-constrained environments has underscored the need for efficient transmission of intermediate feature representations. In this context, feature coding, which compresses features into compact bitstreams, becomes a critical component for scenarios involving feature transmission, storage, and reuse. However, this compression process inevitably introduces semantic degradation that is difficult to quantify with traditional metrics. To address this, we formalize the research problem of Compressed Feature Quality Assessment (CFQA), aiming to evaluate the semantic fidelity of compressed features. To advance CFQA research, we propose the first benchmark dataset, comprising 300 original features and 12000 compressed features derived from three vision tasks and four feature codecs. Task-specific performance degradation is provided as true semantic distortion for evaluating CFQA metrics. We systematically assess three widely used metrics – MSE, cosine similarity, and Centered Kernel Alignment (CKA) – in terms of their ability to capture semantic degradation. Our findings demonstrate the representativeness of the proposed dataset while underscoring the need for more sophisticated metrics capable of measuring semantic distortion in compressed features. This work advances the field by establishing a foundational benchmark and providing a critical resource for the community to explore CFQA. To foster further research, we release the dataset and all associated source code at https://github.com/chansongoal/Compressed-Feature-Quality-Assessment.
Compressed Feature Quality Assessment (CFQA), Coding for Machines, Feature Coding
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3758309isbn: 979-8-4007-2035-2/2025/10ccs: Information systems Multimedia databasesccs: Computing methodologies Image compressionccs: Human-centered computing Ubiquitous and mobile computing design and evaluation methods
@@ 1. Introduction
Rapid deployment of large foundation models (e.g., DINOv2 (Oquab et al., 2023), LLaMA3 (Dubey et al., 2024)) in distributed and resource-constrained environments has created a growing need to transmit intermediate features rather than raw signals (Gao et al., 2024c). Feature coding, which compresses these intermediate representations, plays a vital role in enabling scalable, privacy-preserving, and efficient systems. Unlike traditional image or video coding that prioritizes perceptual quality, feature coding targets the preservation of task-relevant semantics embedded in feature representations.
@@
However, feature coding inevitably introduces semantic degradation: a loss of semantic information that may compromise the performance of large models. This degradation is fundamentally different from pixel-level distortions and often cannot be captured by conventional distortion metrics such as MSE or SSIM. Although task accuracy is a more reliable measurement of semantic distortion, it is impractical for practical deployments: downstream tasks may be inaccessible, costly to run, or unavailable in the feature coding process. These limitations highlight an urgent and underexplored problem: Compressed Feature Quality Assessment (CFQA) – How can we estimate the semantic distortion of compressed features without relying on downstream inference?
@@
Solving CFQA presents several challenges. First, there is no public dataset that provides compressed features with corresponding task performance across multiple tasks and codecs. Second, existing similarity metrics lack validation for high-dimensional features and often fail to generalize across various tasks. Third, the field lacks a unified evaluation protocol to assess whether a given metric reliably measures semantic distortion.
@@
To bridge this gap, we take the first step towards a systematic study of CFQA. In contrast to previous works that rely on task supervision or task-specific codecs, we treat CFQA as a standalone problem and aim to benchmark its core components: dataset, metrics, and evaluation protocols. In summary, we make the following contributions:
@@
  • Benchmark Dataset: We construct the first comprehensive CFQA dataset, consisting of 300 original features and 12000 compressed features from three vision tasks (image classification (Cls), semantic segmentation (Seg), and depth estimation (Dpt)) and four representative codecs (handcrafted and learning-based). The dataset enables quantitative analysis of semantic distortion across tasks, bitrates, and coding methods.
@@
  • Ground-truth Semantic Distortion: For each compressed feature, we provide task-specific semantic distortion labels computed by comparing task head outputs using original and compressed features. These labels serve as the ground truth for training and evaluating semantic distortion metrics.
@@
  • Baseline Metric Evaluation: We evaluate three representative metrics (MSE, cosine similarity, and Centered Kernel Alignment (CKA)) and analyze their sensitivity to feature coding and their ability to measure task-specific semantic distortion. Our analysis reveals their varying strengths and limitations, providing insight into their applicability to semantic distortion measurement.
@@
By establishing a standardized benchmark and evaluation framework, this work bridges the gap between low-level compression and high-level semantic utility. We hope to catalyze the development of lightweight, generalizable, and task-agnostic CFQA metrics that advance the real-world adoption of feature coding. Our dataset and code are publicly available to foster reproducibility and future research.
@@ 2.1. Feature Coding
Feature coding has received growing attention (Zhang et al., 2021; Kim et al., 2023; Misra et al., 2022; Alvar and Bajić, 2019, 2020; Guo et al., 2023; WG2, 2023; Yan et al., 2021; Choi and Bajić, 2018; Chen et al., 2020b, a; Gao et al., 2024a; Zhu et al., 2024; Gao et al., 2025a; Ma et al., 2024; Chen et al., 2024) in edge-cloud collaborative intelligence scenarios. Recently, large model feature coding has attracted increasing interest (Gao et al., 2024c, 2025b, 2025c). These works extend feature coding to scenarios where features are transmitted, stored, and reused.
@@ 2.2. Semantic Distortion Measurement
Task Source Num. of Org. Feat. Pre-processing Feature Codecs Num. of Comp. Feat. Feature Shape GT Distortion
Cls ImageNet 100 Resize
HM, VTM
Multi-task Hyperprior
Task-specific Hyperprior
[…]
@@ 3.1. Problem Formulation
Given a pretrained model $\mathcal{M}$, we denote its extracted intermediate feature from an input $x\in\mathcal{X}$ as $\mathbf{f}=\mathcal{M}(x)\in\mathbb{R}^{d}$, where $d$ is the feature dimension. A feature codec $\mathcal{C}$ encodes $\mathbf{f}$ into a compact bitstream and decodes it back to $\hat{\mathbf{f}}=\mathcal{C}^{-1}(\mathcal{C}(\mathbf{f}))$. Our objective is to evaluate how much semantic information is preserved in $\hat{\mathbf{f}}$ with respect to $\mathbf{f}$, without accessing downstream task ground-truth or executing inference.
[…]
@@ 3.2. Application Scenarios
We illustrate two typical application scenarios of CFQA in Figure 1. In the first scenario (left of Figure 1), CFQA can be directly integrated into the codec training process as a supervisory signal to align compression objectives with downstream task performance. Recent studies (Singh et al., 2020; Gao et al., 2024a) show that task-aware feature coding benefits significantly from semantic-aware supervision. CFQA offers such guidance without requiring end-to-end task labels. As shown on the left of Figure 1, both the original and reconstructed features are input to the CFQA module, which estimates the semantic distortion. The distortion score is then used to generate gradients for optimizing the encoder and decoder. With a well-designed CFQA metric, the resulting codec learns to preserve semantic information even under low-bitrate constraints.
In the second scenario (right of Figure 1), CFQA is essential in edge-cloud collaborative systems, where features are extracted at the edge and transmitted to the cloud for inference. Since the downstream model resides on the cloud side, semantic distortion cannot be directly measured at the edge. Here, CFQA is used as a proxy to estimate task-relevant semantic distortion. As illustrated on the right of Figure 1, the compressed feature is first evaluated by the CFQA module. If it is judged to be of high quality, the bitstream is transmitted to the cloud. Otherwise, the edge device re-encodes the feature at a higher bitrate to better preserve semantic information. This strategy ensures the reliability of transmitted features while reducing unnecessary bandwidth consumption.
Although our work focuses on semantic distortion due to compression, we emphasize that the value of CFQA extends beyond codec benchmarking: it serves as a crucial component in a variety of systems where features are extracted, compressed, transmitted, cached, or reused.
@@ 4.1. Overview
Our proposed dataset is designed to support the analysis of semantic distortion introduced by lossy compression. It includes three tasks, 300 original features, and 12000 compressed features from 4 feature codecs. The features cover diverse image processing methods and multiple feature extraction strategies. The overall information of the dataset is presented in Table 1.
@@ 4.2. Model and Task Selection
Since current feature coding research primarily focuses on visual signals, we initiate the study of CFQA with visual features. We adopt DINOv2 (Oquab et al., 2023) as our backbone feature extractor due to its strong generalization capability and widespread adoption in general-purpose vision tasks. To cover a broad range of semantic distortion, we select three widely studied vision tasks: image classification (Cls), semantic segmentation (Seg), and depth estimation (Dpt). These tasks span coarse-to-fine semantic understanding: Cls focuses on image-level category prediction, Seg introduces spatial semantics with class-wise alignment at the pixel level, and Dpt requires detailed geometry prediction.
This multi-task approach is critical for evaluating the generalizability and sensitivity of CFQA metrics, ensuring they remain robust across different semantic requirements.
@@ 4.3. Source Data Collection
To ensure semantic diversity, we select 100 representative samples for each task. For Cls, we sample 100 ImageNet (Deng et al., 2009) images from 100 distinct categories, each correctly predicted by the DINOv2 classifier. For Seg, we sample 100 images from the Pascal VOC 2012 (Everingham et al., 2010) validation set, covering all 20 semantic categories. For Dpt, we sample 100 images from the NYUv2 (Nathan Silberman and Fergus, 2012) dataset, spanning all 16 scenes. This sampling strategy balances feature diversity with manageable dataset scale, enabling rigorous yet tractable evaluation.
@@ 4.4. Original Feature Collection
The inclusion of diverse image pre-processing techniques and varying split points emulates real-world input variability, enabling rigorous evaluation of CFQA metric generalizability.
@@ 4.5. Compressed Feature Collection
To simulate different types and strengths of semantic distortion, we compress original features through four codecs. All original features are flattened into 2D arrays before encoding.
@@
We select two handcrafted codecs: HM Intra coding (configured with encoder_intra_main_rext.cfg) and VTM Intra coding (configured with encoder_intra_vtm.cfg). Before encoding, we first uniformly quantize the original feature values to [0, 1023]. For both codecs, we use the YUV 4:0:0 (monochrome) format for feature coding and set the quantization parameters to $\in\{2,4,6,8,\ldots,16,18,20\}$ to simulate various bitrates (measured by Bits Per Feature Point, BPFP) and distortion levels.
@@
We adopt the Hyperprior model (Ballé et al., 2018) as it represents a milestone in learning-based feature coding. Since most learning-based feature coding methods build upon this architecture, it serves as an ideal testbed for studying the semantic distortion characteristics of learning-based methods.
To investigate how optimization strategies affect compressed feature quality, we implement two distinct variants. Multi-Task Hyperprior Codec: The codec is trained on features extracted from the three tasks. Task-Specific Hyperprior Codec: The codec is trained exclusively on features extracted from a single task. All codecs are optimized following the training protocols established in (Gao et al., 2024c). The rate-distortion trade-off parameters ($\lambda$) can be found in our released models.
4.6. Semantic Distortion Collection
To establish a rigorous benchmark for evaluating CFQA metrics, we require ground-truth measurements of semantic distortion. We define the true semantic distortion as the performance degradation in downstream tasks when using compressed features $\hat{\mathbf{f}}$ versus original features $\mathbf{f}$. For Cls, we measure the deviation in prediction confidence by computing the rank in the softmax function generated from $\hat{\mathbf{f}}$. The original feature $\mathbf{f}$ achieves perfect ranking (rank=1), with higher ranks indicating more severe semantic degradation. For Seg, we compute the mIoU difference between the segmentation masks predicted from $\hat{\mathbf{f}}$ and $\mathbf{f}$. For Dpt, we compute the RMSE difference between the depth maps predicted from $\hat{\mathbf{f}}$ and $\mathbf{f}$.
These task-specific scores provide quantitative measures of semantic distortion, serving as the foundation for assessing CFQA metric performance.
@@ 4.6. Semantic Distortion Collection
Multi-Task Hyperprior Task-Specific Hyperprior
@@
Hyperprior
(Multi-Task)
Cls -0.0281 0.4707 -0.8907 -0.7165 -0.3277 -0.6593
@@
Hyperprior
(Task-Specific)
Cls -0.0220 0.1303 -0.5486 -0.6084 -0.2622 -0.3365